# Mind the Time: Temporally-Controlled Multi-Event Video Generation

### v1

@@ Abstract
Real-world videos consist of sequences of events. Generating such sequences with precise temporal control is infeasible with existing video generators that rely on a single paragraph of text as input. When tasked with generating multiple events described using a single prompt, such methods often ignore some of the events or fail to arrange them in the correct order. To address this limitation, we present MinT, a multi-event video generator with temporal control. Our key insight is to bind each event to a specific period in the generated video, which allows the model to focus on one event at a time. To enable time-aware interactions between event captions and video tokens, we design a time-based positional encoding method, dubbed ReRoPE. This encoding helps to guide the cross-attention operation. By fine-tuning a pre-trained video diffusion transformer on temporally grounded data, our approach produces coherent videos with smoothly connected events. For the first time in the literature, our model offers control over the timing of events in generated videos. Extensive experiments demonstrate that MinT outperforms existing open-source models by a large margin. Additional results and details are available at our project page.
@@ 1 Introduction
Recent research in video diffusion models [37] has achieved tremendous progress [8, 9, 36, 69, 13, 14, 26, 96]. These approaches typically rely on a single text prompt, and generate videos capturing only a single event. In contrast, real-world videos often comprise sequences of events with rich dynamics. Thus, achieving realism requires the ability to generate multiple events with fine-grained temporal control [63, 88].
Figure 2: Multi-event video generation results from state-of-the-art video generators and MinT. We run two open-source models CogVideoX-5B [100] and Mochi 1 [82], and two commercial models Kling 1.5 [2] and Gen-3 Alpha [1] with a text prompt containing four consecutive events. All of them only generate a subset of events while ignoring the remaining ones. In contrast, MinT generates a natural video with all events smoothly connected. Please refer to Sec. C.6 and our project page for more comparisons.
@@ 3.1 Background: Text-to-Video Latent DiT
where $A_{n,m}$ decreases monotonically with $\lvert n-m\rvert$ when $(n-m)\theta_{\mathrm{base}}\in[-\pi/2,\pi/2]$. This usually holds in our DiT since the video tokens $\bm{z}$ are of low resolution. Please refer to Sec. A.2 for a rigorous discussion. In our video DiT, RoPE is only applied in the self-attention. There is no positional encoding in the video-text cross-attention as the input text prompt is expected to describe the entire video.
@@ 3.2 Temporally Aware Video DiT
As we show in Sec. A.2, our ReRoPE design achieves all three desired properties in the temporal cross-attention.
@@ 4.1 Experimental Setup
Figure 6: Qualitative results of T2V. Concatenating all events into a single prompt (Concat) can only generate the first event. Autoregressive generation (AutoReg) suffers from video stagnation and cannot generate the third event. MEVG struggles to preserve the person’s identity and produces abrupt event transitions. MinT is the only method that generates all events with smooth transitions and consistent content. See Sec. C.1 for more qualitative results and our project page for video results.
@@ 4.2 Text-to-Video Generation
As for multi-event generation methods, AutoReg and MEVG greatly improve the text alignment, as they generate each event from its prompt separately. Yet, AutoReg has much lower visual quality, since conditioning on generated frames leads to artifacts such as video stagnation. MEVG resolves this issue with frame inversion. However, it often generates abrupt transitions between events as indicated by the large number of cuts. In fact, we found that the inversion technique in MEVG only works well when two consecutive event captions have a similar structure, e.g., the same subject doing different actions. When two captions have a subject change such as in Fig. 6, the generated events usually contain completely different characters. Overall, MinT achieves the best balance between video quality, event localization, and temporal smoothness of the video. See Sec. C.6 for comparisons with commercial models.
@@
Human evaluation. We conduct a user study using 200 randomly sampled prompts from HoldOut. We ask the participants to express their preference when presented with paired samples from MinT and each baseline, gathering votes from 5 users per sample. Results in Fig. 7 show that MinT has better or competitive visual quality, while generating events with significantly higher text alignment, timing accuracy, and transition smoothness.
Event time control. MinT supports fine-grained control of event timing. Please refer to Sec. C.4 for our results.
@@ 4.4 Prompt Enhanced Video Generation
Prompt enhancer. We prompt GPT-4 [3] to extend the short prompt to a detailed global caption and temporal captions. Please refer to Sec. C.2 for the prompt we use.
@@
Results. Tab. 2 demonstrates the video generation results from enhanced prompts on VBench. Global performs consistently better than Short, proving that using a detailed prompt is indeed beneficial. When equipped with additional temporal captions, MinT achieves competitive results with baselines in visual quality and motion smoothness, while scoring significantly higher in dynamic degree. Fig. 8 shows a qualitative example, where we turn a single-action prompt into a coherent video with three actions. Please refer to Sec. C.2 for more qualitative results.
@@ 4.5 Ablation Study
ReRoPE rescaling length $L$. By default, we set $L=8$. Tab. 3 shows that using $L=4$ or $16$ achieves similar results. This indicates that the model is insensitive to this hyper-parameter. Please refer to Sec. A.3 for more discussions on ReRoPE with different values of $L$.
Scene cut conditioning. In the last row of Tab. 3, we remove the scene cut conditioning during training. As discussed in Sec. 3.3, without access to scene cut information, the model will introduce undesired shot transitions in the generated video. Indeed, this variant has similar visual quality and text alignment as our full model, but scores much lower in temporal consistency and generates more cuts. Please refer to Sec. C.3 for more analysis.
@@ C.6 Comparison with SOTA Video Generators
C.7 Out-of-Distribution Prompts
@@ Appendix D Limitations and Future Works
Future works. It is interesting to enhance our model with recent progress in training-free long video generation techniques [88, 31, 70]. Another direction is to combine MinT with video personalization methods [40, 50, 54] to enable both fine-grained control within a shot and subject consistency across shots for minute-long video creation.

### v2

@@ Abstract
Real-world videos consist of sequences of events. Generating such sequences with precise temporal control is infeasible with existing video generators that rely on a single paragraph of text as input. When tasked with generating multiple events described using a single prompt, such methods often ignore some of the events or fail to arrange them in the correct order. To address this limitation, we present MinT, a multi-event video generator with temporal control. Our key insight is to bind each event to a specific period in the generated video, which allows the model to focus on one event at a time. To enable time-aware interactions between event captions and video tokens, we design a time-based positional encoding method, dubbed ReRoPE. This encoding helps to guide the cross-attention operation. By fine-tuning a pre-trained video diffusion transformer on temporally grounded data, our approach produces coherent videos with smoothly connected events. For the first time in the literature, our model offers control over the timing of events in generated videos. Extensive experiments demonstrate that MinT outperforms existing commercial and open-source models by a large margin. Additional results and details are available at our project page.
@@ 1 Introduction
Recent research in video diffusion models [39] has achieved tremendous progress [9, 10, 38, 71, 14, 15, 28]. These approaches typically rely on a single text prompt, and generate videos capturing only a single event. In contrast, real-world videos often comprise sequences of events with rich dynamics. Thus, achieving realism requires the ability to generate multiple events with fine-grained temporal control [65, 90].
Figure 2: Multi-event video generation results from SOTA video generators and MinT. We run two open-source models CogVideoX-5B [102] and Mochi 1 [84], and two commercial models Kling 1.5 [2] and Gen-3 Alpha [1] to generate sequential events. All of them only generate a subset of events while ignoring the remaining ones. In contrast, MinT generates a natural video with all events smoothly connected. Please refer to Appendix C.6 and our project page for more comparisons. Comparisons with Sora [12] can be found here.
@@ 3.1 Background: Text-to-Video Latent DiT
where $A_{n,m}$ decreases monotonically with $\lvert n-m\rvert$ when $(n-m)\theta_{\mathrm{base}}\in[-\pi/2,\pi/2]$. This usually holds in our DiT since the video tokens $\bm{z}$ are of low resolution. Please refer to Appendix A.2 for a rigorous discussion. In our video DiT, RoPE is only applied in the self-attention. There is no positional encoding in the video-text cross-attention as the input text prompt is expected to describe the entire video.
@@ 3.2 Temporally Aware Video DiT
As we show in Appendix A.2, our ReRoPE design achieves all three desired properties in the temporal cross-attention.
@@ 4.1 Experimental Setup
Figure 6: Qualitative results of T2V. Concatenating all events into a single prompt (Concat) can only generate the first event. Autoregressive generation (AutoReg) suffers from video stagnation and cannot generate the third event. MEVG struggles to preserve the person’s identity and produces abrupt event transitions. MinT is the only method that generates all events with smooth transitions and consistent content. See Appendix C.1 for more qualitative results and our project page for video results.
@@ 4.2 Text-to-Video Generation
As for multi-event generation methods, AutoReg and MEVG greatly improve the text alignment, as they generate each event from its prompt separately. Yet, AutoReg has much lower visual quality, since conditioning on generated frames leads to artifacts such as video stagnation. MEVG resolves this issue with frame inversion. However, it often generates abrupt transitions between events as indicated by the large number of cuts. In fact, we found that the inversion technique in MEVG only works well when two consecutive event captions have a similar structure, e.g., the same subject doing different actions. When two captions have a subject change such as in Fig. 6, the generated events usually contain completely different characters. Overall, MinT achieves the best balance between video quality, event localization, and temporal smoothness of the video. See Appendix C.6 for comparisons with commercial models.
@@
Human evaluation. We conduct a user study using 200 randomly sampled prompts from HoldOut. We ask the participants to express their preference when presented with paired samples from MinT and each baseline, gathering votes from 20 users per sample. Results in Fig. 7 show that MinT has better or competitive visual quality, while generating events with significantly higher text alignment, timing accuracy, and transition smoothness.
Event time control. MinT supports fine-grained control of event timing. Please refer to Appendix C.4 for our results.
@@ 4.4 Prompt Enhanced Video Generation
Prompt enhancer. We prompt GPT-4 [3] to extend the short prompt to a detailed global caption and temporal captions. Please refer to Appendix C.2 for the prompt we use.
@@
Results. Tab. 2 demonstrates the video generation results from enhanced prompts on VBench. Global performs consistently better than Short, proving that using a detailed prompt is indeed beneficial. When equipped with additional temporal captions, MinT achieves competitive results with baselines in visual quality and motion smoothness, while scoring significantly higher in dynamic degree. Fig. 8 shows a qualitative example, where we turn a single-action prompt into a coherent video with three actions. Please refer to Appendix C.2 for more qualitative results.
@@ 4.5 Ablation Study
ReRoPE rescaling length $L$. By default, we set $L=8$. Tab. 3 shows that using $L=4$ or $16$ achieves similar results. This indicates that the model is insensitive to this hyper-parameter. Please refer to Appendix A.3 for more discussions on ReRoPE with different values of $L$.
Scene cut conditioning. In the last row of Tab. 3, we remove the scene cut conditioning during training. As discussed in Sec. 3.3, without access to scene cut information, the model will introduce undesired shot transitions in the generated video. Indeed, this variant has similar visual quality and text alignment as our full model, but scores much lower in temporal consistency and generates more cuts. Please refer to Appendix C.3 for more analysis.
@@ C.6 Comparison with SOTA Video Generators
Figure 16: Generated videos with extreme dynamics. We prompt MinT to generate scene cuts at event boundaries, leading to explicit scene changes and large dynamics.
@@
Quantitative comparison with Kling. Running Kling on all test prompts will incur unreasonable API costs ($\sim$\$5k). Therefore, we ran it on the 200 prompts used in our user study, and conducted a user study with 20 participants per prompt similar to our main experiment. Due to a weaker base model, MinT achieves a lower Visual Quality (31.55% win rate). Nevertheless, MinT clearly outperforms Kling in all three event-related metrics (73.18% in Text Alignment, 69.93% in Event Timing, and 68.27% in Event Transition).
C.7 Generating Videos with Extreme Dynamics
We prompt MinT to generate videos with extreme dynamics. Thanks to the scene cut conditioning, we enable explicit scene changes in generated videos as shown in Fig. 16.
C.8 Out-of-Distribution Prompts
@@ Appendix D Limitations and Future Works
Future works. It is interesting to enhance our model with recent progress in training-free long video generation techniques [90, 33, 72]. Another direction is to combine MinT with video personalization methods [42, 52, 56, 20] to enable both fine-grained control within a shot and subject consistency across shots for minute-long video creation.